OpenStudy (anonymous):

A moon orbiting a planet in the solar system has an orbital period of 5.0 x 104 s, and the distance between the center of the moon and the planet is 2.0 x 108 m. What planet is it?

OpenStudy (anonymous):

In this question you're aiming to find the mass of this planet so you can use the centripetal force equation and equates it with the gravitational force \[Fc = m w ^{2} r \] and \[F _{G}= G \frac{ m _{planet} m _{Moon} }{ r ^{2} }\] where \[G=6.67×10−11\] and you can find \[w =\frac{ 2\pi }{ T }\] then you can solve it and find the planet's mass.